# COMPACT: COMPositional Atomic-to-Complex Visual Capability Tuning

### v1

@@ 4 Experiments
In this section, we evaluate the baseline approaches and COMPACT on existing multimodal benchmarks. First, we discuss our evaluation setup and the benchmarks (§4.1). Second, we compare the performance of COMPACT with relevant baselines, including LLaVA-665K [24] VIT (§4.2). Third, we analyze COMPACT trained model’s generalization to various levels of compositional complexity and investigate how the distribution of $k$ in the training dataset impacts performance (§4.3). Finally, we conduct ablation studies to understand different aspects of COMPACT design, such as the distribution of compositional complexities, coverage of atomic capabilities, the range of compositional complexities, and the ratio between instruction and compositional tuning data (§4.4).
@@ 4.1 Evaluation Testbed
Model.  We train LLaVA-v1.5-7B-LoRA [24] model’s pre-visual-instruction-tuning checkpoint222 LLaVA-v1.5-mlp2x-336px-pretrain-vicuna-7b-v1.5, which has no prior exposure to visual instruction tuning data. on our COMPACT training dataset. This checkpoint has not been exposed to any visual instruction tuning data prior to COMPACT training. The training dataset includes 32K-sample compositional tuning data unless otherwise stated. Additionally, we mix 5% of LLaVA-665K [24] to preserve instruction following capability. We train the model for one epoch with its official LLaVA-v1.5 LoRA fine-tuning settings.
@@
We compare the effectiveness of our COMPACT data recipe with several baseline datasets by training models with the same architecture under identical training configurations. LLaVA-665K: The full LLaVA-665K [24] VIT dataset (665K samples) used in LLaVA-v1.5. This serves as our primary performance baseline. Random: A 65K-sample random subset of LLaVA-665K [24] that matches our training data size. This baseline controls for data volume. ICONS [39]: A 65K-sample subset of LLaVA-665K [24] selected using the ICONS method, which is a gradient-driven influence-consensus based data selection method that selects the most informative samples for data-efficient visual instruction tuning.
Benchmarks.  We evaluate models trained with different data recipes on established multimodal benchmarks that assess complex visual capabilities. 1) MM-Vet [43] includes 16 types of complex multimodal tasks integrated from 6 core capabilities (recognition, OCR, knowledge, language generation, spatial awareness, and math). 2) MME [9] contains 10 perception (e.g., color, count, OCR) and 4 cognition (e.g., commonsense reasoning, text translation, code understanding) related visual subtasks. 3) LLaVA-in-the-Wild [24] is an open-ended visual question answering benchmark that asks complex questions on real-world images. 4) SeedBench2Plus [17] evaluates visual comprehension skills of MLLMs with a focus on charts, maps, and webs. 5) MMStar [6] contains 1,500 visual questions that span 6 core capabilities (fine-grained perception, coarse perception, mathematics, science & technology, logical reasoning and instance reasoning), carefully curated to evaluate multimodal understanding. 6) CV-Bench [37] is a MLLM benchmark specialized for 2D and 3D visual understanding that includes spatial relationship, object count, relative distance, and depth order. 7) TextVQA [35] evaluates visual understanding of texts in the image. 8) InfoVQA [27] measures visual understanding of infographic images. These benchmarks cover a broad range of vision-centric capabilities. We also note that some of these benchmarks include non-visual questions involving skills as knowledge and math, which are not our primary focus. We provide a more detailed discussion of model performance in these knowledge-intensive and math-intensive tasks in Appendix §A.
@@ 4.2 Main Results
Figure 3: Performance Across Compositional Tuning Data Scales. We fix the VIT subset (5% of LLaVA-665K [24]) and scale the compositional tuning data in COMPACT from 2K to 32K. For comparison, we remove the compositional tuning data and add more VIT data (2K-32K) instead to prepare VIT only recipes with equal data budgets. COMPACT (solid lines) consistently outperforms LLaVA-665K [24] VIT (dashed lines) with fewer data. The performance gap is pronounced for complex reasoning benchmarks such as MM-Vet and MMStar, where the 8K COMPACT model often exceeds the LLaVA-665K [24] VIT baseline at 32K. This demonstrates the data efficiency of COMPACT, requiring substantially less data than LLaVA-665K [24] VIT to achieve comparable or better results.
@@
Overall Performance.  As shown in Tab. 2, COMPACT performs on par with the LLaVA-665K [24] baseline. COMPACT’s training dataset, which amounts to just 10% of the full VIT data, contains a mixture of 32K generated compositional tuning data and 5% of LLaVA-665K [24] VIT data (33K). The compositional tuning data trains the model on compositional capabilities, and the VIT subset maintains the model’s instruction-following capability. COMPACT outperforms both the random [24] and ICONS [39] baselines on most benchmarks, demonstrating superior performance on multi-capability tasks.
@@
COMPACT shows consistent improvements on different benchmarks, achieving strong gains on tasks like MM-Vet [43] (+8.6% over LLaVA-665K [24]), MMStar [6] (+2.9%), InfoVQA [27] (+13.8%), and SeedBench2Plus [17] (+3.4%) while maintaining competitive performance on TextVQA [35] and LLaVA-in-the-Wild [24]. Across all benchmarks, our COMPACT achieves an average relative performance of 100.18%, outperforming even the full LLaVA-665K [24]. In comparison, the random baseline achieves 95.38%, and the ICONS [39] baseline 97.47%, highlighting the effectiveness of our compositional data generation strategy. Additionally, we provide qualitative results in Appendix §C.
@@
Visual Compositional Tuning is Data-Efficient.  We study the data efficiency of COMPACT by analyzing how its performance changes as we scale the amount of compositional tuning data. We fix the VIT subset (5% of LLaVA-665K [24]) and scale the compositional tuning data in COMPACT from 2K to 32K. As comparison, we remove the compositional tuning data and add more VIT data (2K-32K) instead to match the dataset size. Fig. 3 shows that as the number of compositional tuning samples increases, COMPACT performance trends upward across all benchmarks while the random baseline shows mixed behavior as the size of the dataset increases. Furthermore, across all dataset sizes, COMPACT performs consistently better than the random baseline, and the gap increases as the size of the dataset grows. Models trained on smaller compositional tuning data (2K-8K samples) often match or exceed the performance of random baseline models trained on much larger data. For instance, COMPACT’s 2K model achieves 30.73 on MM-Vet [43], outperforming the random baseline’s 32K model at 30.46 which has eight times more data. This demonstrates that COMPACT makes more effective use of training data compared to the baselines.
We hypothesize that this improvement in data efficiency comes from two factors: (1) Balanced Compositional Complexity: COMPACT continuously provides learning signals of higher compositional complexity by balancing the distribution of $k$ in the training dataset. In contrast, the LLaVA-665K [24] VIT paradigm trains almost exclusively on $k=1$ tasks (e.g., single-capability queries). Models trained on the VIT data receive signals of higher compositional complexity less frequently, leaving them unprepared for compositional generalization, the ability to integrate combinations of capabilities not explicitly seen during training (more analysis in §4.3). (2) Diverse Capability Compositions: COMPACT sustains the learning potential during training by explicitly introducing diverse ($k>=1$) integrations of atomic capabilities. Meanwhile, the LLaVA-665K [24] VIT relies heavily on simpler tasks ($k=1$) that can be easily memorized and templatized for rapid saturation of learning potential.
@@ 4.3 Analysis
Figure 4: Compositional Generalization to Higher-Complexities. Performance comparison across compositional complexities ($k$). COMPACT shows competitive performance against LLaVA-665K [24] VIT training. It exceeds the LLaVA-665K [24] baseline at higher compositional complexity tasks ($k=4$ and $k=5$) while using significantly less training data. The $k$-distribution rows show the distribution of compositional complexities in each benchmark.
Performance Gains on Complex Compositional Questions.  COMPACT’s notable performance improvements on complex compositional questions demonstrate its potential for strong compositional generalization. As shown in Fig. 4, COMPACT achieves competitive performance on the MM-Vet [43] and MMStar [6] benchmarks across various levels of compositional complexity ($k$). Despite not being explicitly trained on $k>3$ data, our model effectively generalizes to even higher $k$ regimes. For MM-Vet [43], the scores are 57.5 (COMPACT) vs 32.5 (LLaVA-665K [24]) when $k=4$, and 20.0 (COMPACT) vs 0.0 (LLaVA-665K [24]) when $k=5$. For MMStar [6], the scores are 64.7 (COMPACT) vs 35.3 (LLaVA-665K [24]) when $k=4$. This shows that COMPACT performs robustly in scenarios with higher compositional complexity.
Example Questions with Different Compositional Complexities MM-Vet ($k=3$):
Q: What is the color of the hat worn by the person in the front left?
Required capabilities: color attribution, object recognition, spatial relationship
MMStar ($k=4$):
Q: What is the position of the red rug in the living room?
Required capabilities: color attribution, object recognition, spatial relationship, scene understanding
MMStar ($k=5$):
Q: Is the number of metal cars that are left of the tiny matte school bus greater than the number of tiny cyan double bus?
Required capabilities: spatial relationship, object recognition, counting, color attribution, shape attribution
Distribution of Visual Capabilities.  We use Gemini-2.0-Flash [36] to analyze each question and identify the atomic capabilities required to give an answer (see the details of the system prompt in Appendix §B). Fig. 10 shows the approximate distribution of the number of capabilities required per question in the LLaVA-665K [24] VIT dataset. We sampled 5,668 questions that belong to 1,000 random data points in the VIT dataset, and analyzed their compositional complexity using Gemini-2.0-Flash [36]. The mean compositional complexity of the questions is approximately $k=1.5$, and the mode is $k=1$. 59.2% of the questions utilize only one capability, and an additional 30.9% use 2 capabilities. Together, about 90% of the questions require 2 or less visual atomic capabilities. This complexity cliff in the LLaVA-665K [24] VIT dataset characterized by the scarcity of higher $k$ questions leads to steep declines in its downstream performance on higher $k$ tasks. (Fig. 4). The performance gap between COMPACT and the VIT for different $k$ values shows that the VIT’s complexity-agnostic training leaves models unprepared for tasks that require compositional generalization.
Interestingly, a small fraction of the questions (0.2%) require as many as 10 capabilities (e.g., “Question: Describe this photo in detail.”). We also observe that 1.1% of the questions require zero capabilities, as illustrated in Fig. 10. We further provide $k=0$ examples in Appendix §C. Fig. 5 shows the relative frequencies of atomic capabilities in the question samples. Object recognition (38.97%) and scene understanding (28.58%) are the most common. Other notable capabilities include spatial relationship (25.14%), text recognition (24.68%), and color attribution (14.40%). Less frequent capabilities include object interaction (6.55%), action recognition (6.05%), counting (2.95%), shape attribution (1.13%), and spatial recognition (1.06%).
Figure 5: Comparison of Capability Distribution. The heatmaps show the frequency of each atomic capability in LLaVA (left) and COMPACT (right) samples. The capabilities are sorted by frequency based on the LLaVA capability distribution, with more common capabilities appearing closer to the top. In LLaVA, the distribution is notably imbalanced: object recognition and scene understanding are some of the most frequent, while shape and spatial recognition are less prevalent. In contrast, our COMPACT exhibits a more balanced distribution across all capability categories.
Recipe #Data InfoVQA [27] SeedBench2Plus [17] MME [9] TextVQA [35] MMVet [43] CV-Bench [37] MMStar [6] LLaVA-W [24] Rel. (%)
LLaVA-665K [24] 665K 20.80 41.72 1478.48 46.99 29.22 60.92 35.11 68.50 100.00
Random 49K 20.33 42.38 1290.45 42.22 30.18 54.75 34.3 70.5 96.28
Unbalanced COMPACT 49K 22.28 41.17 1339.24 43.08 29.22 55.84 34.8 64.5 96.62
COMPACT 49K 22.68 42.82 1362.68 43.73 30.78 54.69 35.59 66.6 98.83
Table 3: Matching LLaVA-665K Distribution. Performance comparison of unbalanced COMPACT and multiple baselines. The distribution of compositional complexity in unbalanced COMPACT follows LLaVA-665K [24]. Training a model on unbalanced COMPACT leads to performance on par with training on the random baseline which is a subset of LLaVA-665K [24] equal in size, suggesting that a balanced distribution of $k$ in compositional tuning data is critical in compositional generalization.
@@ 4.4 Ablation Studies
Figure 6: Leave-One-Out Analysis on Atomic Capabilities. We measure the average performance degradation across benchmarks by excluding an atomic capability from training. Higher drop indicates higher importance of the atomic capability. Excluding scene understanding and spatial relationships have the largest impact, while that of excluding shape and action recognition are modest.
We conduct a series of ablation studies to investigate key design considerations (compositional complexity distribution, atomic capability coverage, compositional complexity range, and instruction tuning ratio) in COMPACT. Unless otherwise specified, all experiments use 5% of LLaVA-665K [24] VIT data and 16K $k=1,2,3$ compositional tuning data.
Effect of Matching LLaVA-665K Distribution.  In order to show that the performance improvement of COMPACT mainly comes from the balanced distribution of compositional complexity in the compositional tuning data, we analyze the impact on performance when its compositional complexity is unbalanced. We further generate a 16K-sample compositional tuning data whose distribution of $k$ resembles that of LLaVA-665K [24], which is heavily skewed as in Fig. 10. This gives us 58,168 $k=1$, 30,364 $k=2$, and 7,468 $k=3$ conversations. Similar to the original COMPACT data recipe, we mix the unbalanced 16K compositional tuning data with the random 5% subset of the VIT data. We compare this unbalanced COMPACT training dataset with the following baselines: 1) a same size random subset of the VIT data, 2) the original COMPACT with 16K balanced compositional tuning data, and 3) the full VIT dataset. As shown in Tab. 3, the performance of unbalanced COMPACT stands at 96.62% (in relation to the full baseline), close to the random baseline at 96.28%. However, the performance of original COMPACT jumps to 98.83%, suggesting that most of the performance gain in COMPACT comes from the fair representation of higher $k$ samples in the compositional tuning data.
Impact of Atomic Capability Coverage.  To validate our choice of atomic capabilities and understand their relative importance, we conduct a leave-one-out analysis by systematically excluding questions that require a specific capability while keeping the total number of training examples fixed. As shown in Fig. 6, scene understanding and spatial relationship emerge as the most critical capabilities, with each of their exclusion leading to a significant performance drop (5.22% and 4.93% respectively). Text recognition and object recognition are also essential (4.65% and 4.03% drops). The exclusion of capabilities like shape attribution and action recognition have a smaller impact (0.74% and 2. 08%). This analysis validates our selection of atomic capabilities by demonstrating that each capability contributes meaningfully to overall performance without being redundant.
Figure 7: Compositional Complexity Analysis: Performance comparison of models trained with different compositional complexities. $k=1$ refers to only one atomic capability per question, $k=1,2$ to both single and dual capabilities, and $k=1,2,3$ to single, dual, and triple capabilities. Results show consistent improvements as the range of compositional complexities increases.
Effect of Compositional Complexity Range.  To isolate the effect of the range of compositional complexities while controlling for data quality, we generate three sets of 16K-sample compositional tuning data, each with $k=1$, $k=1,2$ or $k=1,2,3$, using identical Gemini-2.0-Flash [36] configurations. For fair comparison, we maintain consistent sample counts and use an identical set of images in all three settings. The model trained on only $k=1$ (single capability per question) underperforms the model trained on $k=1,2,3$ compositional tuning data on multi-capability benchmarks: MM-Vet [43] (28.82 vs. 29.22), LLaVA-W [24] (66.1 vs. 68.5) and MMStar [6] (34.53 vs. 35.11). This shows that although the model trained on $k=1$ data can solve tasks with lower compositional complexity, it struggles to perform in higher compositional complexities.
As shown in Fig. 7, increasing the range of compositional complexities leads to consistent improvements on all three benchmarks. Training on $k=1,2,3$ compositional tuning data achieves the highest performance on MM-Vet [43] (32.61) and MMStar [6] (0.3577), demonstrating that exposure to more complex compositional patterns during training enhances the model’s ability to handle complex multi-capability tasks. Surprisingly, the model achieves 112% performance on MM-Vet [43] with only 16k compositional tuning data compared to the LLaVA-665K [24] baseline, suggesting that a balanced mixture of different compositional complexities improves data efficiency.
Impact of Instruction Tuning Data Ratio.  We vary the amount of instruction tuning data sampled from the LLaVA-665K [24] VIT data to understand the impact of the mixing ratio on model performance. In order to isolate the effect on visual instruction following, we exclude $k=0$ conversations (approximately 1.1% of the questions), which have minimal relevance to visual capabilities. As we scale the VIT subset from 0% (pure compositional tuning) to 7% of LLaVA-665K [24], we observe an upward trend in performance, indicating that the role of the instruction tuning data is crucial. Fig. 8 shows that without instruction tuning data (0%), COMPACT achieves only 74.69% of the performance relative to LLaVA-665K [24] VIT. Increasing the instruction tuning data to just 1% of the VIT data significantly improves the relative performance to 96.56%. However, further scaling gives diminishing returns, with 3% reaching 98.77% and 5% achieving nearly identical relative performance (99.99%). Interestingly, taking 7% from the VIT data causes a slight decrease to 98.07%, indicating that 5% represents an optimal balance between instruction tuning and compositional tuning data in terms of data efficiency and performance. These results suggest that instruction following capability is potentially orthogonal to the capabilities of the base model and the atomic visual capabilities, and can be acquired with minimal instruction tuning data.
@@
Figure 8: Impact of Instruction Tuning Data Ratio on Performance. Relative performance of models trained on COMPACT mixed with different ratios instruction tuning data from LLaVA-665K [24]. The x-axis is the percentage of LLaVA-665K [24] used as instruction tuning data, and the y-axis is the average relative score across benchmarks. The performance improves significantly with a small percentage of instruction tuning data and stabilizes around 5%.

### v2

@@ 4 Experiments
In this section, we evaluate the baseline approaches and COMPACT on existing multimodal benchmarks. First, we discuss our evaluation setup and the benchmarks (§4.1). Second, we compare the performance of COMPACT with relevant baselines, including LLaVA-665K (llava15) VIT (§4.2). Third, we investigate the source of COMPACT’s performance improvement. (§4.3). Finally, we ablate the various design choices of COMPACT4.4).
@@ 4.1 Evaluation Testbed
Model.  We train LLaVA-v1.5-7B-LoRA (llava15) model’s pre-visual-instruction-tuning checkpoint [^2] LLaVA-v1.5-mlp2x-336px-pretrain-vicuna-7b-v1.5, which has no prior exposure to visual instruction tuning data. on our COMPACT training dataset. This checkpoint has not been exposed to any visual instruction tuning data prior to COMPACT training. The training dataset includes 32K-sample compositional tuning data unless otherwise stated. Additionally, we mix 5% of LLaVA-665K (llava15) to preserve instruction following capability. We train the model for one epoch with its official LLaVA-v1.5 LoRA fine-tuning settings.
@@
We compare the effectiveness of our COMPACT data recipe with several baseline datasets by training models with the same architecture under identical training configurations. LLaVA-665K: The full LLaVA-665K (llava15) VIT dataset (665K samples) used in LLaVA-v1.5. This serves as our primary performance baseline. Random: A 65K-sample random subset of LLaVA-665K that matches our training data size. This baseline controls for data volume. We further evaluate on 65K-sample subsets curated with various data reduction methods: EL2N (el2n), Perplexity (perplexity), SemDeDup (semdedup), D2-Pruning (d2), Self-Sup (selfsup), Self-Filter (self-filter), and ICONS (wu2024icons).
Benchmarks.  We evaluate models trained with different data recipes on established multimodal benchmarks that assess complex visual capabilities. 1) MM-Vet (mmvet) includes 16 types of complex multimodal tasks integrated from 6 core capabilities (recognition, OCR, knowledge, language generation, spatial awareness, and math). 2) MME (mme) contains 10 perception (e.g., color, count, OCR) and 4 cognition (e.g., commonsense reasoning, text translation, code understanding) related visual subtasks. 3) LLaVA-in-the-Wild (llava15) is an open-ended visual question answering benchmark that asks complex questions on real-world images. 4) SeedBench2Plus (seedbench) evaluates visual comprehension skills of MLLMs with a focus on charts, maps, and webs. 5) MMStar (mmstar) contains 1,500 visual questions that span 6 core capabilities (fine-grained perception, coarse perception, mathematics, science & technology, logical reasoning and instance reasoning), carefully curated to evaluate multimodal understanding. 6) CV-Bench (cvbench) is a MLLM benchmark specialized for 2D and 3D visual understanding that includes spatial relationship, object count, relative distance, and depth order. 7) TextVQA (textvqa) evaluates visual understanding of texts in the image. 8) InfoVQA (infovqa) measures visual understanding of infographic images. These benchmarks cover a broad range of vision-centric capabilities. We also note that some of these benchmarks include non-visual questions involving such skills as knowledge and math, which are not our primary focus. We provide a more detailed discussion of these tasks in Appendix §A.2 and  §A.3.
@@ 4.2 Main Results
Overall Performance.  As shown in Tab. 2, COMPACT performs on par with the LLaVA-665K (llava15) baseline with only 10% of its data volume. COMPACT’s training dataset contains a mixture of 32K compositional tuning data and 5% of LLaVA-665K (llava15) VIT data (33K). The compositional tuning data trains the model on complex questions, and the VIT subset maintains the model’s instruction-following capability.
@@
Across all benchmarks, our COMPACT achieves an average relative performance of 100.2%, outperforming the full LLaVA-665K (llava15) (100.0%), random baseline (95.4%), and ICONS (wu2024icons) (97.5%). COMPACT achieves strong gains on tasks like MM-Vet (mmvet) (+8.6% over LLaVA-665K (llava15)), MMStar (mmstar) (+2.9%), InfoVQA (infovqa) (+13.8%), and SeedBench2Plus (seedbench) (+3.4%) while maintaining competitive performance on TextVQA (textvqa) and LLaVA-in-the-Wild (llava15). This highlights the effectiveness of the COMPACT data recipe. Additionally, we provide qualitative results in Appendix §C.
@@
Complex Training Data is Efficient.  We study the data efficiency of complex samples by analyzing how COMPACT’s performance changes as we scale the amount of compositional tuning data. We fix the VIT subset (5% of LLaVA-665K (llava15)) and scale the compositional tuning data in COMPACT from 2K to 32K. For comparison, we match the size of each dataset purely with VIT data. Fig. 3 shows that as the number of compositional tuning data samples increases, the performance of COMPACT on multimodal benchmarks steadily improves unlike the random baseline. Interestingly, models trained on smaller amount of compositional tuning data (2K-8K samples) often match or exceed the performance of random baseline models trained on much larger VIT data. For instance, COMPACT’s 2K model achieves 30.8 on MM-Vet (mmvet), outperforming the random baseline’s 32K model at 30.5. This demonstrates that COMPACT makes more effective use of training data compared to the baselines.
This improvement in data efficiency comes from several factors. COMPACT training data has higher $k$ samples that encourage the model to extract more visual information from the image. The learning potential of lower $k$ samples that LLaVA-665K (llava15) relies on are limited by the amount of visual information that the model can afford to ignore. However, a range of simple to complex training samples are still necessary to properly disentangle each atomic capability from complex questions. COMPACT’s complexity-aware data generation process that progressively increases the complexity of the samples achieves balance these effects.
Table 3: Matching LLaVA-665K distribution. We show that performance improvements of COMPACT can be attributed to increased complexity $k$. $\textsc{COMPACT}_{llava}$ capability distribution exactly matches that of random LLaVA-665K subset, controlling for complexity. Note that we use 16K compositional data here (different from 32K in Tab. 2) for faster iteration. Roughly half of the performance gain in COMPACT can be attributed to the higher $k$ compositional tuning data.
Recipe #Data InfoVQA SeedBench2Plus MME TextVQA MMVet CV-Bench MMStar LLaVA-W Rel. (%)
LLaVA-665K 665K 20.80 41.72 1478.48 46.99 29.22 60.92 35.11 68.50 100.00
Random 49k 20.33 42.38 1290.45 42.22 30.18 54.75 34.3 70.5 96.28
$\textsc{COMPACT}_{llava}$ 49k 22.76 43.43 1308.52 43.94 28.81 52.39 36.08 66.8 97.55
COMPACT 49k 22.68 42.82 1362.68 43.73 30.78 54.69 35.59 66.6 98.83
@@ 4.3 Analysis
Complexity Distribution in LLaVA-665K We characterize the complexity distribution of LLaVA-665K (llava15) VIT dataset that serves as the primary baseline for COMPACT. We use Gemini-2.0-Flash (team2023gemini) to analyze the $k$-value of 5,400 questions that belong to 1,000 randomly sampled images (see the details of the system prompt in Appendix §B). Fig. 1 shows that the mean $k$ value of the samples is approximately $k=1.95$, and the mode is $k=2$. 77% of the questions require two or fewer atomic visual capabilities. Interestingly, a small fraction of the questions (0.06%) require as many as 10 capabilities (e.g., “Question: Describe this photo in detail.”). We also observe that 0.9% of the questions require zero capabilities. We further provide $k=0$ examples (see Appendix §C.2).
Complexity Distribution in COMPACT We confirm that COMPACT increases the complexity distribution of the training dataset previously established by LLaVA-665K. We use Gemini-2.0-Flash (team2023gemini) to analyze 7200 questions that belong to 1000 randomly sampled images in the compositional tuning data. The mean $k$-value of the data is $k=2.89$, and mode is $k=3$, which are higher than the LLaVA counterparts. We also find that the representation of each capability in the dataset is more balanced compared to LLaVA-665K (see Appendix Fig. 6). We provide more details on the data statistics of COMPACT and LLaVA-665K data in Appendix §A.5.
COMPACT successfully scales the $k$-value distribution of the training dataset. The number of capabilities sampled during generation ($k_{gen}$) in the Capability Sampling step (§3.2) practically acts as the lower bound of the question’s final $k$-value for two reasons: 1) Object recognition capability is implicitly assumed by the conversation generator (i.e. Gemini (team2023gemini)) during generation and verification steps (see Fig. 8 in Appendix). We hypothesize that “object-oriented”ness is implicitly required to generate high-quality questions about the image, likely due to the biases in the image sources. We note that COMPACT does not change the visual contents of the images. 2) We find that spatial capabilities such as scene understanding, spatial recognition, and spatial relationship are often related in practice as the question becomes spatially complex (see Fig. 8 in Appendix). The spatial orientation of objects can be influenced by the overall layout of the scene. For example, what is “on the left side of” an object could be described as “behind”, depending on their relative depth.
Effect of Matching LLaVA-665K Distribution.  We isolate the complexity effect from COMPACT’s performance gain by controlling for the sample generator. We generate a 16K-sample compositional tuning data whose capability and $k$-value distribution matches that of LLaVA-665K (llava15). Similar to the original COMPACT data recipe, we mix this complexity-matched compositional tuning data with the random 5% subset of the VIT data. We compare this training dataset with the following baselines: 1) a random subset of the VIT data with the same size, 2) original COMPACT with 16K compositional tuning data, and 3) the full VIT dataset. As shown in Tab. 3, the performance of complexity-matched COMPACT ($\textsc{COMPACT}_{llava}$) stands at 97.6% (compared to the full VIT baseline). The performance of original COMPACT jumps to 98.8%, suggesting that at least half of the performance gain (1.3% out of 2.6% improvement over random baseline) in COMPACT comes from higher complexity in the compositional tuning data.
@@ 4.4 Ablation Studies
Figure 4: Impact of k-value range: Performance comparison across variations of COMPACT whose compositional tuning data is synthesized with different ranges of $k_{gen}$ in the Capability Sampling step (§3.2). Models trained on higher and wider ranges of complexity (darker bars) achieve higher performance across benchmarks.
@@
Figure 5: Impact of instruction tuning data ratio. Relative performance of COMPACT with different amounts instruction tuning data from LLaVA-665K (llava15). The x-axis is the percentage of LLaVA-665K used as instruction tuning data, and the y-axis is relative score. The performance improves significantly with a small amount of instruction tuning data and stabilizes around 5%.
We conduct ablation studies to understand how COMPACT addresses the aforementioned challenges in using complexity as tool to control sampling for VIT data curation. First, we vary the range of complexity in the data to show that training efficiency benefits from higher and wider ranges of $k$. Second, we show that response formatting can be trained by mixing a relatively small amount of prompt-engineered VIT dataset. Unless otherwise specified, all experiments use 5% of LLaVA-665K (llava15) VIT data and 16K compositional tuning data.
Effect of k-value Range.  We analyze the relationship between information density and efficient learning based on the learning potential of different ranges of complexity. We train the model on COMPACT with five different versions of 16K compositional tuning data, each generated from using $k_{gen}=1$, $k_{gen}=2$, $k_{gen}=3$, $k_{gen}=1,2$ or $k_{gen}=1,2,3$ in the Capability Sampling step (§3.2). For fair comparison, we maintain consistent sample counts and use an identical set of images in all five settings. As shown in Fig. 5, increasing the number of sampled capabilities per question leads to consistent improvements on all three benchmarks. Training on $k_{gen}=1,2,3$ compositional tuning data achieves the highest performance on MM-Vet (mmvet) and LLaVA-in-the-wild (llava15), and second highest on MMStar (mmstar). Surprisingly, sampling from $k_{gen}=1,2,3$ demonstrates stronger performance on average compared to sampling only $k_{gen}=3$. This suggests that although complex samples are more information dense, their benefit is maximized in the presence of simpler samples. We provide more details on k-value range in Appendix §A.2.
Impact of Instruction Tuning Data Ratio.  We show that the domain adaptation problem of fitting model’s response to the format required by the test dataset can be resolved with a data mixture. We fix the 16K compositional tuning data and scale the amount of VIT subset from 0% (pure compositional tuning) to 7% of LLaVA-665K (llava15). Fig. 5 shows that a small addition of VIT subset improves the performance of the model on MMVet (mmvet) (short answer), SeedBench2Plus  (seedbench) (multiple choice) and LLaVA-in-the-Wild  (llava15) (long answer). We find that the prompt engineered VIT subset is training the instruction following capability of the model on various question types even at just 1%. Interestingly, further scaling gives diminishing returns, and starts to lose absolute gains at 7%. These results suggest that instruction following capability is potentially orthogonal to the capabilities of the base model and the atomic visual capabilities, and can be acquired with minimal instruction tuning data.